Select the functions that are equivalent to A(t) = 5,000 (0.976 )^t , Where t is the time in years.

Group of answer choices
A(t) = 5,000 ( 0. 72 -t )^t
A(t) = 5,000 (0.947)^6t
A(t) = 5,000 (9.24)^t
A(t) = 5,000 ( 0.998)^12t
A(t) = 5,000 ( 0. 976 1/6 )^6t

Answer by ikleyn(52781)

.
Select the functions that are equivalent to A(t) = 5,000 (0.976 )^t , Where t is the time in years.

Group of answer choices
A(t) = 5,000 ( 0. 72 -t )^t
A(t) = 5,000 (0.947)^6t
A(t) = 5,000 (9.24)^t
A(t) = 5,000 ( 0.998)^12t
A(t) = 5,000 ( 0. 976 1/6 )^6t
~~~~~~~~~~~~~~~


No one of listed functions is equivalent.


Do you mock the tutors ?


ONLY  if to write the last function properly,  i.e. in the form

            A(t) = 5,000 ( (0. 976)^(1/6) )^(6t),

then it will be equivalent.